A bomb of mass 1 kg is thrown vertically upwards with a speed of 100 m/s. After 5 seconds it explodes into 2 fragments. One fragment of mass 400 gm is found to go down with a speed of 25 m/s. What will happen to the second fragment just after the explosion? (g = 10m/s)

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Velocity of the bomb after 5 seconds(at the time of explosion)=u-gt=100-(10)(5)=50m/supwards

Since no external force acts on the system in the explosion, the linear momentum of the system remains conserved.

Hencem_{bomb}v_{bomb}=m_1v_1+m_2v_2
\implies (1)(+50)=(0.4)(-25)+(0.6)v_2
\implies v_2=+100m/s
Hence the other fragment will move upwards with 100m/s.